Assume the Knight Corporation is considering the acquisition of Day Inc. The expected earnings per share for the Knight Corporation will be $10 with or without the merger. However, the standard deviation of the earnings will go from $1.80 to $1.50 with the merger because the two firms are negatively correlated.


a. Compute the coefficient of variation for the Knight Corporation before and after the merger. (Do not round intermediate calculations and round your answers to 2 decimal places.)
  


b. Comment on the possible impact on Knight’s postmerger P/E ratio, assuming investors are risk-averse.
Answer #1

1.
Before merger=1.80/10=0.18
After merger=1.50/10=0.15

2.
As risk decreases required return would decrease hence price would increase hence PE ratio might increase. People would be willing to pay more as risk decreases.